OpenStudy (anonymous):

Which of the following is the correct classification of the system of equations below? -x+2y=10 6y=-12x+1 Parallel Coincident Intersecting

OpenStudy (campbell_st):

put them both on point slope form 2y = x + 10 or y = 1/2x + 5 slope m = 1/2 6y = -12x + 1 or y = -2x + 1/6 slope m = -2 the product of the gradients is 1/2 x -2 -1 so the lines are perpendicular

OpenStudy (anonymous):

so Which one is it ?

OpenStudy (anonymous):

so Which one is it ?

OpenStudy (campbell_st):

they intersect at right angles.... that is the definition of perpendicular lines

OpenStudy (anonymous):

oh ok thanks